# Critical Batch Size Revisited: A Simple Empirical Approach to Large-Batch Language Model Training

### v2

@@ 4.1 Methodology for Batch Size Warmup
Given an existing training run (at a small batch size), we aim to adapt it to achieve the the same final loss with fewer overall gradient steps (i.e., a larger batch size for most of training). Assuming an original fixed batch size $B$ and base learning rate $\eta$, our method for training with batch size warmup is as follows. First, we use branched training CBS method (Section 3) to measure $B^{*}_{t}$, the CBS after training for $t$ tokens, using an existing checkpoint. When training a new model, we set the batch size and and learning rate as follows:
Batch Size Warmup 1. At $t=0$, initialize the batch size to $B_{0}=B$ and base learning rate to $\eta_{0}=\eta$. 2. After training for $t$ tokens, if we determine that the CBS exceeds the current batch size ($B^{*}_{t}>2B_{t}$), we double the current batch size and update the base learning rate following the square-root scaling rule (cf. Section 3): $B_{t+1}$ $=2B_{t}$ $\eta_{t+1}$ $=\sqrt{2}\eta_{t}.$
@@
The main implementation question is how to operationalize checking $B^{*}_{t}>2B_{t}$ in order to double the batch size. In practice, we do this heuristically based on the measurements in Figure 2: since we only double the batch size twice over training, this involves just picking two thresholds. In addition, this process could be automated for future training runs using online measurements of the CBS paired with some kind of curve fitting or statistical test.
@@
As defined above, our method uses a square-root scaling rule, which is well-motivated for Adam (Malladi et al., 2022), but in principle a linear scaling rule could also be used. As batch size warmup only modifies the base learning rate, it is compatible to overlay with existing learning rate schedules (in practice, the models we use follow a cosine schedule). Finally, we choose to only increase the batch size at powers of two because it is not clear that having a more precise match to the CBS beyond the order-of-magnitude level would be particularly useful and because the OLMo codebase requires the number of GPUs divides the batch size. Thus, it is convenient to only increase the batch size when we know the new batch size would remain an integer. However, in principle it would be possible to adapt our method to update the batch size by multipliers $\gamma<2$.
@@
Our batch size warmup method is similar to preliminary experiments by McCandlish et al. (2018, Appendix D) using a dynamic batch size on the SVHN dataset. However, they use their noise scale method (Section 2) to set the batch size, which is potentially unreliable, and only consider small-scale classification. In the context of image classification, prior work explored replacing learning rate decay with increasing the batch size (Smith et al., 2018). This is conceptually related to our batch size warmup approach in that it leverages scaling rules between batch size and learning rate to achieve larger batch training. However, our method can apply on top of existing learning rate schedules, can generalize to Adam (vs. SGD), and, most crucially, ensures that the batch size number exceeds the critical batch size.
@@ 4.2 Experimental Setup
    Batch Size Warmup: We train OLMo 1B with our batch size warmup method as detailed in Section 4.1. We initialize the batch size to 1024. Based on a manual reading of the CBS measurements in Section 3, we determine that the CBS reaches 2048 by 168B tokens and 4096 by 503B tokens. Our method thus doubles the batch size at each of these points.
@@
[…]
@@
Language model training consists of two steps, pre-training and mid-training (OLMo et al., 2025). We evaluate training loss as well as several measures of out-of-distribution loss at the end of the pre-training stage as well as after the mid-training stage.
@@
  • Loss after Mid-Training. We report loss at the end of mid-training stage, which more closely reflects how these checkpoints are used in language model training. Specifically, starting from final pretraining checkpoint, we linearly anneal the remaining learning rate down to 0 for 50B tokens, keeping the batch size fixed at its final value (Figure 1). Because mid-training is a part of the standard language modeling pipeline, we take the loss after mid-training to represent the loss that would be achieved by these training runs in a practical context. Furthermore, OLMo et al. (2025) suggest that this kind of learning rate annealing can induce significant gains in loss for partial pretraining runs. We thus also take the loss after mid-training loss as a proxy for how these runs would compare if we were to fully train them for the full learning rate schedule.
  • Out-of-Distribution Losses. To measure performance beyond the training loss, we also track three kinds of out-of-distribution losses. The first two are straightforward cross-entropy losses on common pretraining datasets, C4 (Dodge et al., 2021) and The Pile (Gao et al., 2020). The third out-of-distribution loss follows the method from Bhagia et al. (2024), which argues for computing the loss (in bits-per-byte; BPB) on the correct answers of multiple popular question-answering datasets such as ARC-Easy, ARC-Challenge, MMLU, etc. For the full list of datasets used, see Appendix E.
@@ 4.3 Batch Size Warmup Results
Table 2: According to the method from Bhagia et al. (2024), we evaluate downstream performance via BPB on downstream tasks, as well as cross-entropy loss on two held-out sets, C4 and the Pile, both after pretraining and after mid-training. Batch-size warmup generally performs comparably or better compared to the small-batch control, suggesting it does not degrade downstream performance.
We compare batch size warmup to the small and large-batch controls at two points: first, after the pretraining stage and, second, after the mid-training stage that anneals the remaining learning rate down to 0. After pretraining, batch size warmup reaches lower loss than both controls, as shown in Figure 2 and Table 1, outperforming the small-batch control in loss by a margin of 0.0166. After mid-training, batch size warmup still slightly outperforms the small-batch control in loss, this time by a smaller margin of 0.0053. The large-batch control now achieves the worst overall loss. Overall, batch size warmup slightly exceeds the small-batch control in loss while using 43% fewer gradient steps. In contrast, the large-batch control uses 75% fewer gradient steps, but its final loss degrades compared to the small-batch control. Thus, we find batch size warmup is a reliable method to train with fewer gradient steps without degrading final loss.
Beyond the final loss, we also evaluate whether batch size warmup leads to similar downstream performance as small-batch training using our downstream-task BPB evaluation as well as validation loss on C4 and the Pile. As shown in Table 2, we find that BPB on validation sets is broadly competitive with the small-batch control, performing best on three out of four conditions considered. With the caveat that precisely measuring the downstream impact of pretraining decisions is difficult, this suggests that, just as batch-size warmup does not degrade final loss, it should not degrade downstream measures of performance either.
@@ 5 Conclusion
In this work, we introduced a simple empirical method for estimating the CBS throughout language model pretraining runs, which be used to increase the batch size (and thus effective token throughput) for large scale training runs without sacrificing performance. As we discussed, the existing noise scale method (McCandlish et al., 2018) for estimating the CBS requires strong assumptions to justify, which our approach avoids. We used our method to study the evolution of the CBS during training for the OLMo models, finding that CBS increases monotonically but diminishing over the course of training, and that CBS does not seem to depend on model size, in line with prior work (Zhang et al., 2024). Guided by these findings, we showed that our measurements could be used to pick a batch size warmup schedule that enables larger batch training without harming final training loss. We take these results to demonstrate the validity and utility of our CBS measurement approach, and believe our framework could be useful for increasing the efficiency of future large-scale pretraining efforts.
@@ Appendix A CBS Measurement Details
Our empirical method for measuring the CBS Section 3 is, in principle, sensitive to the choice of checkpoints and batch size multipliers. We therefore document the checkpoints and multipliers we used here.
@@
When measuring the OLMo 1B CBS with branching, we set the base batch size to 1024 and the base learning rate to $0.0004\cdot\sqrt{2}$, reflecting the default batch size of 512 and learning rate of 0.0004 in the OLMo codebase under a square-root scaling rule [Malladi et al., 2022]. We then chose the following checkpoints and multipliers $k$:
@@
Figure 5 shows loss vs. batch size plots for all checkpoints of OLMo 1B.
@@
Appendix A shows loss vs. batch size plots for all checkpoints of OLMo 7B. These checkpoints were chosen manually as we developed this project. Over the course of our experimentation, we launched many additional runs beyond the ones discussed above. Since the choice of $k$ can influence the conclusions of our method, we filtered down the included runs to make the choice of $k$ systematic.
@@ Appendix B Noise Scale Measurement Details
We use the gradient noise scale estimator proposed by McCandlish et al. [2018, Appendix A] to estimate the gradient noise scale. The method estimates the gradient noise scale using gradient norms at two different batch sizes $B_{\mathrm{big}}$ and $B_{\mathrm{small}}$ according to:
@@
We estimate a confidence interval for $\mathcal{B}_{\mathrm{simple}}$ in two steps. First, we estimate 95% confidence intervals for $\mathcal{S}$ and $\lVert\mathcal{G}\rVert^{2}$, assuming the data are exponentially [^2] For the exponential distribution, we use approximate confidence interval under “Confidence Intervals” here: https://en.wikipedia.org/wiki/Exponential_distribution. and normally distributed, respectfully, based on manual inspection of their distributions (cf. Figure 6). We denote these intervals $[a_{\mathcal{S}},b_{\mathcal{S}}]$ and $[a_{\lVert\mathcal{G}\rVert^{2}},b_{\lVert\mathcal{G}\rVert^{2}}]$, respectively. We then define the confidence interval for $\mathcal{B}_{\mathrm{simple}}$ as follows:
@@ Appendix C License Information
The OLMo models [Groeneveld et al., 2024, OLMo et al., 2025] and pretraining code, which we use, are released under Apache-2.0 license. C4 [Dodge et al., 2021] is released under ODC-BY license. The Pile [Gao et al., 2020] is released under MIT license.
@@ Appendix D Deriving CBS Scaling Laws: An Attempt
To begin, we assume that the goal of picking a fixed batch size $B$ is to minimize the L2 distance to the local CBS over the course of training. It is not obvious that minimizing L2 distance is the right way to pick the fixed CBS: for instance, we might want to weight training at a batch size above the local CBS more negatively than training below it. Regardless, we will proceed for now under the assumption that this is the right perspective. We also make the weaker assumption that $f(t)=0$, in line with our empirical findings (Section 3). It follows that the best batch size to train at (i.e., fixed CBS) is simply the average local CBS over training:
@@
Note that the second derivative $2T$ is positive. Thus, setting the derivative to 00 and solving for $B$, we conclude that the following value of $B$ minimizes $R_{2}$:
@@
Thus, under the assumptions we have made, if we are trying to pick a fixed batch size that best approximates the local CBS throughout training, we can simply pick the average CBS over training. We can use Proposition 1 to derive a scaling law for the fixed $B^{*}$ as a function of the final CBS or, equivalently, the total steps $T$. We now consider various reasonable functional forms $f(t)$ for the CBS.
@@ D.1 Power Law CBS Scaling
In the case where $c=1/2$ (square root), $B^{*}=\frac{2}{3}B^{*}_{T}=\frac{2}{3}\sqrt{T}$, which derives the $\sqrt{T}$ scaling law proposed by prior work [Zhang et al., 2024].
@@ Appendix E BPB Evaluation on Downstream Tasks
This section lists the datasets we used to compute BPB measures for downstream tasks. For multiple-choice tasks, we use the Cloze/Completion Formulation (CF), and compute the BPB metric on the gold answer. For completion tasks, we simply compute BPB over the correct answer. This approach was inspired by Bhagia et al. [2024]. The selection of tasks follows the guidelines from Magnusson et al. [2025].

### v3

@@ 4.1 Methodology for Batch Size Warmup
Given an existing training run (at a small batch size), we aim to adapt it to achieve the the same final loss with fewer overall gradient steps (i.e., a larger batch size for most of training). Assuming an original fixed batch size $B$ and base learning rate $\eta$, our method for training with batch size warmup is as follows. First, we use branched training CBS method (Section˜3) to measure $B^{*}_{t}$, the CBS after training for $t$ tokens, using an existing checkpoint. When training a new model, we set the batch size and and learning rate as follows:
Batch Size Warmup 1. At $t=0$, initialize the batch size to $B_{0}=B$ and base learning rate to $\eta_{0}=\eta$. 2. After training for $t$ tokens, if we determine that the CBS exceeds the current batch size ($B^{*}_{t}>2B_{t}$), we double the current batch size and update the base learning rate following the square-root scaling rule (cf. Section˜3): $B_{t+1}$ $=2B_{t}$ $\eta_{t+1}$ $=\sqrt{2}\eta_{t}.$
@@
The main implementation question is how to operationalize checking $B^{*}_{t}>2B_{t}$ in order to double the batch size. In practice, we do this heuristically based on the measurements in Figure˜2: since we only double the batch size twice over training, this involves just picking two thresholds. In addition, this process could be automated for future training runs using online measurements of the CBS paired with some kind of curve fitting or statistical test.
@@
As defined above, our method uses a square-root scaling rule, which is well-motivated for Adam (malladi2022sdes), but in principle a linear scaling rule could also be used. As batch size warmup only modifies the base learning rate, it is compatible to overlay with existing learning rate schedules (in practice, the models we use follow a cosine schedule). Finally, we choose to only increase the batch size at powers of two because it is not clear that having a more precise match to the CBS beyond the order-of-magnitude level would be particularly useful and because the OLMo codebase requires the number of GPUs $g$ divides the batch size. Thus, doubling is convenient because it easily guarantees the new batch size remains a multiple of $g$. However, in principle one could update more frequently, e.g., by setting updating the batch size the largest multiple of $g\leq B^{*}_{t}$.
@@
Our batch size warmup method is similar to preliminary experiments by mccandlish2018empirical using a dynamic batch size on the SVHN dataset. However, they use their noise scale method (Section˜2) to set the batch size, which is potentially unreliable, and only consider small-scale classification. In the context of image classification, prior work explored replacing learning rate decay with increasing the batch size (smith-2018-decay). This is conceptually related to our batch size warmup approach in that it leverages scaling rules between batch size and learning rate to achieve larger batch training. However, our method can apply on top of existing learning rate schedules, can generalize to Adam (vs. SGD), and, most crucially, ensures that the batch size number exceeds the critical batch size.
@@ 4.2 Experimental Setup
    Batch Size Warmup: We train OLMo 1B with our batch size warmup method as detailed in Section˜4.1. We initialize the batch size to 1024. Based on a manual reading of the CBS measurements in Section˜3, we determine that the CBS reaches 2048 by 168B tokens and 4096 by 503B tokens. [^1] These thresholds were determined from preliminary CBS measurements at an earlier stage of the project. Our measurements changed slightly after this point, but we deemed that it was not worth it to restart the expensive training runs because we do not think the results should be sensitive to a precise choice of threshold. Going forward, it could be useful to establish systematic methodology for choosing batch size warmup thresholds given the CBS measurements. Our method thus doubles the batch size at each of these points.
@@
    Small-Batch Control: We train OLMo 1B with batch size $B=1024$ and base learning rate $\eta=\sqrt{2}\cdot 0.0004$. [^2] The default hyperparameters for OLMo 1B use a batch size of 512 and learning rate of 0.0004. Since we chose an initial base size of 1024 for our experiments, we adapted the learning rate appropriately via the square-root scaling rule (malladi2022sdes). We expect that our batch size warmup method should be able to achieve similar final loss to the small-batch control with fewer gradient steps.
@@
Language model training consists of two steps, pre-training and mid-training (olmo20252olmo2furious). We evaluate training loss as well as several measures of out-of-distribution loss at the end of the pre-training stage as well as after the mid-training stage.
@@
  • Loss after Mid-Training. We report loss at the end of mid-training stage, which more closely reflects how these checkpoints are used in language model training. Specifically, starting from final pretraining checkpoint, we linearly anneal the remaining learning rate down to 0 for 50B tokens, keeping the batch size fixed at its final value (Figure˜1). Because mid-training is a part of the standard language modeling pipeline, we take the loss after mid-training to represent the loss that would be achieved by these training runs in a practical context. Furthermore, olmo20252olmo2furious suggest that this kind of learning rate annealing can induce significant gains in loss for partial pretraining runs. We thus also take the loss after mid-training loss as a proxy for how these runs would compare if we were to fully train them for the full learning rate schedule.
  • Out-of-Distribution Losses. To measure performance beyond the training loss, we also track three kinds of out-of-distribution losses. The first two are straightforward cross-entropy losses on common pretraining datasets, C4 (C4) and The Pile (Pile). The third out-of-distribution loss follows the method from Bhagia2024EstablishingTS, which argues for computing the loss (in bits-per-byte; BPB) on the correct answers of multiple popular question-answering datasets such as ARC-Easy, ARC-Challenge, MMLU, etc. For the full list of datasets used, see Appendix E.
@@ 4.3 Batch Size Warmup Results
Table 2: According to the method from Bhagia2024EstablishingTS, we evaluate downstream performance via BPB on downstream tasks, as well as cross-entropy loss on two held-out sets, C4 and the Pile, both after pretraining and after mid-training. Batch-size warmup generally performs comparably or better compared to the small-batch control, suggesting it does not degrade downstream performance.
We compare batch size warmup to the small and large-batch controls at two points: first, after the pretraining stage and, second, after the mid-training stage that anneals the remaining learning rate down to 0. After pretraining, batch size warmup reaches lower loss than both controls, as shown in Figure˜2 and Table˜1, outperforming the small-batch control in loss by a margin of 0.0166. After mid-training, batch size warmup still slightly outperforms the small-batch control in loss, this time by a smaller margin of 0.0053. The large-batch control now achieves the worst overall loss. Overall, batch size warmup slightly exceeds the small-batch control in loss while using 43% fewer gradient steps. In contrast, the large-batch control uses 75% fewer gradient steps, but its final loss degrades compared to the small-batch control. Thus, we find batch size warmup is a reliable method to train with fewer gradient steps without degrading final loss.
Beyond the final loss, we also evaluate whether batch size warmup leads to similar downstream performance as small-batch training using our downstream-task BPB evaluation as well as validation loss on C4 and the Pile. As shown in Table˜2, we find that BPB on validation sets is broadly competitive with the small-batch control, performing best on three out of four conditions considered. With the caveat that precisely measuring the downstream impact of pretraining decisions is difficult, this suggests that, just as batch-size warmup does not degrade final loss, it should not degrade downstream measures of performance either.
@@ 5 Conclusion
In this work, we introduced a simple empirical method for estimating the CBS throughout language model pretraining runs, which can be used to increase batch size (and thus effective token throughput) for large scale training runs without sacrificing performance. As we discussed, the existing noise scale method (mccandlish2018empirical) for estimating the CBS requires strong assumptions to justify, which our approach avoids. We used our method to study the evolution of the CBS during training for the OLMo models, finding that CBS increases monotonically but diminishing over the course of training, and that CBS does not seem to depend on model size, in line with prior work (zhang2024cbs). Guided by these findings, we showed that our measurements could be used to pick a batch size warmup schedule that enables larger batch training without harming final training loss. We take these results to demonstrate the validity and utility of our CBS measurement approach, and believe our framework could be useful for increasing the efficiency of future large-scale pretraining efforts.
There are several details and extensions of our method that would be interesting to explore going forward. First, it would be interesting to carry out a more systematic analysis of the impact of the hyperparameter $\Delta$ on CBS measurements: how sensitive is the method to $\Delta$ and do different values for $\Delta$ potentially bias the measurement? It would also be interesting to further investigate the conditions under which noise scale might be a meaningful CBS proxy and used for batch size warmup. When applying our CBS measurements to picking a batch size warmup schedule, we manually picked doubling threshold based on the CBS. Going forward, it would be useful to establish a systematic way to set threshold for increasing the batch size given CBS measurements. There are also other potential methodological improvements, such as removing the arbitrary power of 2 constraint and estimating the CBS in an online fashion. Overall, these improvements would allow batch size warmup to be applied more robustly and easily across different pretraining setups.
@@ Appendix A CBS Measurement Details
Our empirical method for measuring the CBS Section˜3 is, in principle, sensitive to the choice of checkpoints and batch size multipliers. We therefore document the checkpoints and multipliers we used here.
@@
When measuring the OLMo 1B CBS with branching, we set the base batch size to 1024 and the base learning rate to $0.0004\cdot\sqrt{2}$, reflecting the default batch size of 512 and learning rate of 0.0004 in the OLMo codebase under a square-root scaling rule [malladi2022sdes]. We then chose the following checkpoints and multipliers $k$:
@@
Figure˜5 shows loss vs. batch size plots for all checkpoints of OLMo 1B.
@@
Appendix˜A shows loss vs. batch size plots for all checkpoints of OLMo 7B. These checkpoints were chosen manually as we developed this project. Over the course of our experimentation, we launched many additional runs beyond the ones discussed above. Since the choice of $k$ can influence the conclusions of our method, we filtered down the included runs to make the choice of $k$ systematic.
@@ Appendix B Noise Scale Measurement Details
We use the gradient noise scale estimator proposed by mccandlish2018empirical to estimate the gradient noise scale. The method estimates the gradient noise scale using gradient norms at two different batch sizes $B_{\mathrm{big}}$ and $B_{\mathrm{small}}$ according to:
@@
We estimate a confidence interval for $\mathcal{B}_{\mathrm{simple}}$ in two steps. First, we estimate 95% confidence intervals for $\mathcal{S}$ and $\lVert\mathcal{G}\rVert^{2}$, assuming the data are exponentially [^3] For the exponential distribution, we use approximate confidence interval under “Confidence Intervals” here: https://en.wikipedia.org/wiki/Exponential_distribution. and normally distributed, respectfully, based on manual inspection of their distributions (cf. Figure˜6). We denote these intervals $a_{\mathcal{S}},b_{\mathcal{S}}]$ and $a_{\lVert\mathcal{G}\rVert^{2}},b_{\lVert\mathcal{G}\rVert^{2}}]$, respectively. We then define the confidence interval for $\mathcal{B}_{\mathrm{simple}}$ as follows:
@@ Appendix C License Information
The OLMo models [groeneveld-etal-2024-olmo, olmo20252olmo2furious] and pretraining code, which we use, are released under Apache-2.0 license. C4 [C4] is released under ODC-BY license. The Pile [Pile] is released under MIT license.
@@ Appendix D Deriving CBS Scaling Laws: An Attempt
To begin, we assume that the goal of picking a fixed batch size $B$ is to minimize the L2 distance to the local CBS over the course of training. It is not obvious that minimizing L2 distance is the right way to pick the fixed CBS: for instance, we might want to weight training at a batch size above the local CBS more negatively than training below it. Regardless, we will proceed for now under the assumption that this is the right perspective. We also make the weaker assumption that $f(t)=0$, in line with our empirical findings (Section˜3). It follows that the best batch size to train at (i.e., fixed CBS) is simply the average local CBS over training:
@@
Note that the second derivative $2T$ is positive. Thus, setting the derivative to 0 and solving for $B$, we conclude that the following value of $B$ minimizes $R_{2}$:
@@
Thus, under the assumptions we have made, if we are trying to pick a fixed batch size that best approximates the local CBS throughout training, we can simply pick the average CBS over training. We can use Proposition˜1 to derive a scaling law for the fixed $B^{*}$ as a function of the final CBS or, equivalently, the total steps $T$. We now consider various reasonable functional forms $f(t)$ for the CBS.
@@ D.1 Power Law CBS Scaling
In the case where $c=1/2$ (square root), $B^{*}=\frac{2}{3}B^{*}_{T}=\frac{2}{3}\sqrt{T}$, which derives the $\sqrt{T}$ scaling law proposed by prior work [zhang2024cbs].
@@ Appendix E BPB Evaluation on Downstream Tasks
This section lists the datasets we used to compute BPB measures for downstream tasks. For multiple-choice tasks, we use the Cloze/Completion Formulation (CF), and compute the BPB metric on the gold answer. For completion tasks, we simply compute BPB over the correct answer. This approach was inspired by Bhagia2024EstablishingTS. The selection of tasks follows the guidelines from DataDecide.